# Securing the Skies: A Comprehensive Survey on Anti-UAV Methods, Benchmarking, and Future Directions

### v2

@@ 5 Anti-UAV Tracking Approaches
(1) Detection-Based Trackers. Detection-based pipelines usually prioritize real-time performance and utilize efficient object detectors. For example, Wu et al. [144] proposed EECF, a dual-fovea correlation filter tracker operating at 86 FPS with a state accuracy score of 66.42% on Anti-UAV [58] and 70.45% on USC-Drone [24]. Wang et al. [134] integrated YOLOX [44] with Swin Transformer [77] and ByteTrack, achieving 78.2% accuracy on UAVSwarm [133]. Deng et al. [32] combined YOLOv9 [136] and EfficientNet-B7 [123] with LSTM/Kalman filtering for 3D tracking, earning top-ranked pose MSE on MMAUD [156] and winning the UG2+ Challenge [90].
@@
(2) Siamese (Contrastive) Trackers. Siamese trackers excel in learning robust feature similarities and handling fast or evasive UAV motion. Huang et al. [49] claimed first-place in the 1st and 2nd Anti-UAV Challenges [58, 162] with SiamSTA, which incorporates sophisticated spatio-temporal attention and motion estimation. Li et al. [71] augmented SuperDiMP with a dual-branch IR fusion strategy, boosting precision by 4.52% in the 2nd Anti-UAV Challenge. Chen et al. [22] further enhanced SiamSTA with CDCF and a multi-stage re-detection mechanism, excelling against small, fast UAVs. Beyond single-modality, Zhang et al. [161]’s SiamFusion seamlessly integrated RGB and IR streams for a 63.37 score on Anti-UAV, illustrating the tangible value of complementary sensor modalities.
@@
(3) Hybrid (Detection + Siamese) Trackers. Hybrid methods combine detection pipelines with Siamese matching, capitalizing on each paradigm’s respective strengths. Cheng et al. [25] introduced SiamAD, a synthesis of YOLOv5, multi-hybrid attention modules, and SiamRPN++, achieving 88.4% detection precision at 37.1 FPS on the Anti-UAV dataset [58]. Yu et al. [155] integrated Transformers, YOLOv5, LoFTR, GMM, and OSTrack for second place in the 3rd Anti-UAV Challenge [164], affirming the practical effectiveness of multi-stage IR-based tracking.
@@
Key Insights & Challenges. Although real-time Siamese trackers [111] and detection-based pipelines [40] can sustain high frame rates, they often struggle under clutter, occlusion, or swarm conditions, thereby motivating more adaptive architectures [150, 23]. Stealth maneuvers and jamming further complicate UAV tracking, emphasizing the necessity for robust domain adaptation and dynamic cue prioritization [34]. While RGBT sensor fusion [67, 125] mitigates individual sensor shortcomings, incorporating LiDAR or radar [32] inevitably increases calibration overhead and computational load. Existing benchmarks rarely capture large-scale swarms, diverse weather conditions, or severe occlusions, consequently restricting methodical progress [145]. Transformer-based solutions [23, 30, 146] address complex motion patterns and partial occlusion but can lag in real-time performance, necessitating specialized hardware optimizations for on-board or edge devices. Extending multi-object trackers (e.g., ByteTrack [159]) to larger UAV swarms remains challenging due to association uncertainty and domain-specific constraints. Ultimately, next-generation trackers will demand faster inference, improved sensor fusion, and broader datasets to achieve reliable anti-UAV performance in real-world conditions.
@@ 6 Emerging Techniques & Future Directions
Diffusion Models for Data Synthesis & Domain Adaptation. Diffusion models [48, 116] generate high-fidelity UAV imagery to address data scarcity and enhance domain diversity, including text-conditioned approaches [97, 105]. Recent advances in faster sampling [106, 95] support near real-time augmentation, while diffusion-based domain adaptation [94, 13, 160] narrows the gap between synthetic and real-world data. Future research may focus on on-the-fly adaptation from live sensor streams, enabling continuous refinement under dynamic UAV operating conditions.
@@
Vision-Language Models & Semantic Context. Vision-language (VL) methods integrate textual cues with visual features to facilitate open-vocabulary UAV detection and classification [96, 167, 157, 45]. UAV-specific adaptations [76, 52, 73] incorporate textual descriptions on flight profiles, mission goals, or payload attributes, thereby significantly enriching threat assessment. By leveraging semantic narratives (e.g., infiltration routes, suspect cargo), future VL-based systems can swiftly adapt to novel UAV categories and enhance situational awareness.
@@
Self-Supervised & Unsupervised Learning. Self- and unsupervised approaches alleviate dependence on labeled aerial data across varying altitudes, sensors, and weather [64, 65, 152]. Recent work addresses swarm-wide generalization [84], further strengthening multi-UAV detection [63], tracking [137, 66], and recognition [114], as well as security [4]. Moving forward, large-scale self-supervised pre-training that rapidly adapts to domain shifts in real time may enable ongoing responses to evolving UAV behaviors and stealth tactics.
@@
Multi-Modal Fusion & Sensor Integration. Combining sensors (e.g., RGB, IR, radar, and audio) significantly boosts UAV detection and tracking robustness [92, 82, 107, 171]. Classical fusion [92, 75] is now augmented by deep embedding methods [7, 151], allowing adaptive sensor weighting under real-time constraints. However, maintaining precise synchronization [143] and carefully balancing power or bandwidth remain essential considerations for robust, large-scale deployments.
@@
Reinforcement Learning & Autonomous Control. Reinforcement Learning (RL) has proven effective for adaptive UAV decision-making in dynamic environments, building on foundational breakthroughs [83, 112]. Modern UAV-centric RL research tackles interception [15, 173], detection [6], tracking [142], and high-speed racing [62], frequently integrating generative or multi-modal inputs for rapid policy updates. Future directions may emphasize hierarchical or multi-agent RL for large-scale swarm control, balancing cooperative behavior with adversarial countermeasures and mission-specific constraints.

### v3

@@ 5 Anti-UAV Tracking Approaches
(1) Detection-Based Trackers. Detection-based pipelines usually prioritize real-time performance and utilize efficient object detectors. For example, Wu et al. [Wu2024EagleEye] proposed EECF, a dual-fovea correlation filter tracker operating at 86 FPS with a state accuracy score of 66.42% on Anti-UAV [anti_uav] and 70.45% on USC-Drone [USC_Drone]. Wang et al. [wang2024target] integrated YOLOX [YOLOX] with Swin Transformer [liu2021swin] and ByteTrack, achieving 78.2% accuracy on UAVSwarm [wang2022uavswarm]. Deng et al. [deng2024multi] combined YOLOv9 [Wang2024YOLOv9LW] and EfficientNet-B7 [Tan2019EfficientNetRM] with LSTM/Kalman filtering for 3D tracking, earning top-ranked pose MSE on MMAUD [yuan2024mmaud] and winning the UG2+ Challenge [UG2Challenge2024].
@@
(2) Siamese (Contrastive) Trackers. Siamese trackers excel in learning robust feature similarities and handling fast or evasive UAV motion. Huang et al. [Huang2021SiamSTA] claimed first-place in the 1st and 2nd Anti-UAV Challenges [anti_uav, zhao20212nd] with SiamSTA, which incorporates sophisticated spatio-temporal attention and motion estimation. Li et al. [li2022dual] augmented SuperDiMP with a dual-branch IR fusion strategy, boosting precision by 4.52% in the 2nd Anti-UAV Challenge. Chen et al. [Chen2022SiamSTA] further enhanced SiamSTA with CDCF and a multi-stage re-detection mechanism, excelling against small, fast UAVs. Beyond single-modality, Zhang et al. [zhang2023modality]’s SiamFusion seamlessly integrated RGB and IR streams for a 63.37 score on Anti-UAV, illustrating the tangible value of complementary sensor modalities.
@@
(3) Hybrid (Detection + Siamese) Trackers. Hybrid methods combine detection pipelines with Siamese matching, capitalizing on each paradigm’s respective strengths. Cheng et al. [Cheng2022AntiUAVTracking] introduced SiamAD, a synthesis of YOLOv5, multi-hybrid attention modules, and SiamRPN++, achieving 88.4% detection precision at 37.1 FPS on the Anti-UAV dataset [anti_uav]. Yu et al. [Yu2023UnifiedTransformer] integrated Transformers, YOLOv5, LoFTR, GMM, and OSTrack for second place in the 3rd Anti-UAV Challenge [zhao20233rd], affirming the practical effectiveness of multi-stage IR-based tracking.
@@
Key Insights & Challenges. Although real-time Siamese trackers [shen2022real] and detection-based pipelines [Fang2021Real] can sustain high frame rates, they often struggle under clutter, occlusion, or swarm conditions, thereby motivating more adaptive architectures [yan2021learning, chen2021transformer]. Stealth maneuvers and jamming further complicate UAV tracking, emphasizing the necessity for robust domain adaptation and dynamic cue prioritization [do2025ramots]. While RGBT sensor fusion [Li2019Multi, tang2024revisiting] mitigates individual sensor shortcomings, incorporating LiDAR or radar [deng2024multi] inevitably increases calibration overhead and computational load. Existing benchmarks rarely capture large-scale swarms, diverse weather conditions, or severe occlusions, consequently restricting methodical progress [Wu2021Survey]. Transformer-based solutions [chen2021transformer, cui2022mixformer, xiao2024target] address complex motion patterns and partial occlusion but can lag in real-time performance, necessitating specialized hardware optimizations for on-board or edge devices. Extending multi-object trackers (e.g., ByteTrack [zhang2022bytetrack]) to larger UAV swarms remains challenging due to association uncertainty and domain-specific constraints. Ultimately, next-generation trackers will demand faster inference, improved sensor fusion, and broader datasets to achieve reliable anti-UAV performance in real-world conditions.
@@ 6 Emerging Techniques & Future Directions
Diffusion Models for Data Synthesis & Domain Adaptation. Diffusion models [ho2020denoising, song2021score] generate high-fidelity UAV imagery to address data scarcity and enhance domain diversity, including text-conditioned approaches [ramesh2022dalle2, saharia2022imagen]. Recent advances in faster sampling [salimans2022progressive, qiao2025fast] support near real-time augmentation, while diffusion-based domain adaptation [peng2024unsupervised, benigmim2023one, zhang2025domain] narrows the gap between synthetic and real-world data. Future research may focus on on-the-fly adaptation from live sensor streams, enabling continuous refinement under dynamic UAV operating conditions.
@@
Vision-Language Models & Semantic Context. Vision-language (VL) methods integrate textual cues with visual features to facilitate open-vocabulary UAV detection and classification [Radford2021LearningTV, Zhong_2022_CVPR, Zareian_2021_CVPR, li2024human, dong2025ha, dong2025unified]. UAV-specific adaptations [Liu_2024_ESWA, Huang_2025_Drones, Li_2024_ECCV] incorporate textual descriptions on flight profiles, mission goals, or payload attributes, thereby significantly enriching threat assessment. By leveraging semantic narratives [cheng2024shield] (e.g., infiltration routes, suspect cargo), future VL-based systems can swiftly adapt to novel UAV categories and enhance situational awareness.
@@
Self-Supervised & Unsupervised Learning. Self- and unsupervised approaches alleviate dependence on labeled aerial data across varying altitudes, sensors, and weather [Kim_2019_ICCV, Kim_2019_CVPR, Ye_2022_CVPR]. Recent work addresses swarm-wide generalization [Mou2024luster], further strengthening multi-UAV detection [10473588], tracking [Wang_2021_IJCV, Lai_2020_CVPR], and recognition [Sio_2020_MM], as well as security [Alanazi2024SSRL]. Moving forward, large-scale self-supervised pre-training that rapidly adapts to domain shifts in real time may enable ongoing responses to evolving UAV behaviors and stealth tactics.
@@
Multi-Modal Fusion & Sensor Integration. Combining sensors (e.g., RGB, IR, radar, and audio) significantly boosts UAV detection and tracking robustness [park2015radarAudio, messina2019radar, samaras2019multisensor, anti_uav_600]. Classical fusion [park2015radarAudio, liu2017audioCamera] is now augmented by deep embedding methods [Alla2024, Yang2024AVFDTI], allowing adaptive sensor weighting under real-time constraints. However, maintaining precise synchronization [Wu2024multi] and carefully balancing power or bandwidth remain essential considerations for robust, large-scale deployments.
@@
Reinforcement Learning & Autonomous Control. Reinforcement Learning (RL) has proven effective for adaptive UAV decision-making in dynamic environments, building on foundational breakthroughs [mnih2015human, silver2016mastering]. Modern UAV-centric RL research tackles interception [bertoin2022autonomous, cetin2022], detection [alkhonaini2024], tracking [wisniewski2024towards], and high-speed racing [kaufmann2023champion], frequently integrating generative or multi-modal inputs for rapid policy updates. Future directions may emphasize hierarchical or multi-agent RL for large-scale swarm control, balancing cooperative behavior with adversarial countermeasures and mission-specific constraints.